Three years ago, Mary was seven times as old as her sister Jane was. Two years ago, Mary was four times as old as Jane was, and one year ago she was three times as old as Jane was. How old are Mary and Jane now?
Answer by scott8148(6628)

m-3=7(j-3) ... m=7j-18

m-1=3(j-1) ... (7j-18)-1=3j-3 ... 4j=16 ... j=4

m=7(4)-18 ... m=10